OpenStudy (skarlettrose186):

Given: 2x + 3y = 6. What is the y-intercept? (0, 3) (2, 0) (0, 2) no intercept

jhonyy9 (jhonyy9):

do you can isolate the y term ?

OpenStudy (skarlettrose186):

y=\[\frac{ 2x-6 }{ 3 }\]

OpenStudy (seratul):

What is -6/3?

OpenStudy (skarlettrose186):

-2

OpenStudy (seratul):

Wait, you made a mistake. When isolating the y intercept, you subtract 2x on both sides. 3y=2x-6

OpenStudy (seratul):

I mean 3y=-2x+6

OpenStudy (skarlettrose186):

ok so y=-2x+6/3

OpenStudy (seratul):

Then, you divide by three and get y=-2/3x+3 This is now put into y=mx+b form b is always the y intercept. So 3 would be the y intercept. To find x, plug 3 for y. You would get x =0. Therefore, it would be (0,3)

OpenStudy (skarlettrose186):

ok thanks

OpenStudy (seratul):

No problem.

jhonyy9 (jhonyy9):

Rose do you understand it now ?

jhonyy9 (jhonyy9):

why i ve said try isolate the term of y ?

OpenStudy (skarlettrose186):

yes